OpenStudy (anonymous):

solve 25^x-30(5^x) +125=0 a) x=1,2 b) x=-1,2 c) x=5,25 d)x= -5,25

OpenStudy (anonymous):

ans : a) 5^2x-30(5^x) +125=0 put 5^x=t t^2-30t+125=0 , solve it for t and compaire the powers you will get x=1, 2